OpenStudy (anonymous):

Had a question about dividing polynomials and this is the specific question.... 15x^4-10x^3+11 divided by -15x^4

OpenStudy (anonymous):

I divided everything in the numerator by -15x^4 so it should look like this: 15x^4 / -15x^4 -10x^3 / -15x^4 11/-15x^4

OpenStudy (anonymous):

Yes, you are almost there

OpenStudy (anonymous):

So you got -1+2/(3x)-11/(15x^4)

OpenStudy (anonymous):

so the 15x^4 / -15x^4 cancel out. right? the -10x^3 / -15x^4 is what's getting me.... I get 10x^3 / 15x^4 = 3/4x^-1 = 1/ 3/4x = 1 1/3x? 11/-15x^4 = -11/15x^-3 which turns into... ... \[a^-n = \left(\begin{matrix}1 \\ a^n\end{matrix}\right)\]

OpenStudy (anonymous):

Why woule -10x^3/(-15x^4) give you 3/(4x^(-1))?

OpenStudy (anonymous):

10/15=2/3 and x^3/x^4=1/x=x^(-1)

OpenStudy (anonymous):

so 10x^3/(15x^4)=2/(3x)=(2/3)x^(-1)